OpenStudy (anonymous):

x^4 - 4 = 0 How do you solve?

OpenStudy (anonymous):

factorise using a^2-b^2=(a+b)(a-b)

OpenStudy (anonymous):

whats A and B? 1 and -4?

OpenStudy (anonymous):

x^4 - 4 = 0 (x^2)^2 -2^2=0 (x^2-2)(x^2+2)=0 (x+sqrt(2))(x-sqrt(2))(x+isqrt(2))(x-isqrt(2))=0 hence x=+sqrt(2), -sqrt(2) ,+i*sqrt(2), -i*sqrt(2)

OpenStudy (anonymous):

a---> x^2 and b-->2

OpenStudy (anonymous):

OHH yeah! thanks, i think i remember now